8. The first-order line of 589 nm light falling on a diffraction grating is observed at a 15.5 angle. a) How far apart are the slits? b) At what angle will the third order be observed? [1]
Answer #1


In diffraction


path difference = a*sintheta


a = slit width

for constructive path difference = m*lambda/2


lambda = wavelength = 589 nm

m = order of fringe = 1


a*sintheta = m*lambda/2


a*sin15.5 = 1*589*10^-9/2

a = 1.1*10^-6 m

============


sintheta = m*lambda/(2*a)

sintheta = 3*589*10^-9/(2*1.1*10^-6)


theta = 53.4 degrees <<---answer